OpenStudy (anonymous):

find the coordinates of the y-intercept of the line....y=4x-14

OpenStudy (anonymous):

put x=0 y=4(0)-14 y=14 so coordinates of y intercepts are (0,-14)

OpenStudy (lgbasallote):

that should be y = -14 not y = 14 little typo there lol

OpenStudy (anonymous):

will it be the same for y=-4/3x+3/4

OpenStudy (anonymous):

y=3/4 ?

OpenStudy (anonymous):

@ineedhelponmyhw you are right for y=-4/3x+3/4 it should be 3/4

OpenStudy (anonymous):

thank you!